Suppose that the marginal propensity to consume is

dC
dy

= 0.7 − e−2y     (in billions of dollars)

and that consumption is $5.7 billion when disposable income is $0. Find the national consumption function.
Answer #1

Given = 0.7 -e de = (072e-29). dy fac = f (6.7-629) dy C: 0.79 -e tE l e= Integration constant) -2 C 20079 + ,2te 0 Given (=substitute ein to ear o I The national consumption I function is : a Lc.o.7yte +5.2 CS Scanned with CamScanner